a pay phone is holding its maximum number of 500 coins consisting of nickels dimes and quarters. the number of quarters is twice the numbers of dimes. if the value of all the coins is $88 how many nickels were in the pay phone?

Answer by checkley79(3341)

Q+D+N=500 OR 2D+D+N=500 OR 3D+N=500 0R N=500-3D
Q=2D
.25Q+.10D+.05N=88
.25*2D+.10D+.05(500-3D)=88
.50D+.10D+25-.15D=88
.45D=88-25
.45D=63
D=63/.45
D=140 DIMES.
Q=2*140=280 QUARTERS.
280+140+N=500
N=500-280-140
N=80 NICKELS.
PROOF:
.25*280+.10*140+.05*80=88
70+14+4=88
88=88